The acceleration of gravity at the surface of MArs is -3.7m/s squared. what is the initial velocity of a stone thrown upward by Marvin and reaches a height of 19 meter?
This question might relate to the quadratic equation but I really just feel like im just missing one part of this question to solve it, please help.

Answer by nyc_function(2741)

Step 1.) Find the appropriate kinematic equation: You are given Δy (the height, 19 meters), the acceleration of gravity, and you are trying to solve for the initial velocity. They didn't give you time. I would go with v^2 = v0^2 + 2gΔy.
Step 2.) Plugging the equation in:
v = 0 (it's reasonable to think that the speed at maximum height is 0; the stone goes up at a certain speed, stops, and then falls down)
g = -3.7 (the acceleration is given in the problem)
Δy = 19 (the height is given in the problem)
Find v0 (the initial velocity)
v^2 = v0^2 + 2gΔy
0^2 = v0^2 + 2(-3.7)(19)
0 = v0^2 + (-140.6)
140.6 = v0^2
v0 ≈ 11.857 m/s